A right triangle has a hypotenuse of length 15 and a leg of length 11. What is the area of the triangle? If necessary, round your answer to two decimal places.
Answer by edjones(8007)

a^2+b^2=c^2
11^2+b^2=15
121+b^2=225
b^2=104
b=2sqrt(26)
.
A=.5ab
.5*11*2sqrt(26)
=56.09
.
Ed